### v2

# Neural Fields for Adaptive Photoacoustic Computed Tomography

Abstract
Photoacoustic computed tomography (PACT) is a non-invasive imaging modality with wide medical applications. Conventional PACT image reconstruction algorithms suffer from wavefront distortion caused by the heterogeneous speed of sound (SOS) in tissue, which leads to image degradation. Accounting for these effects improves image quality, but measuring the SOS distribution is experimentally expensive. An alternative approach is to perform joint reconstruction of the initial pressure image and SOS using only the PA signals. Existing joint reconstruction methods come with limitations: high computational cost, inability to directly recover SOS, and reliance on inaccurate simplifying assumptions. Implicit neural representation, or neural fields, is an emerging technique in computer vision to learn an efficient and continuous representation of physical fields with a coordinate-based neural network. In this work, we introduce NF-APACT, an efficient self-supervised framework utilizing neural fields to estimate the SOS in service of an accurate and robust multi-channel deconvolution. Our method removes SOS aberrations an order of magnitude faster and more accurately than existing methods. We demonstrate the success of our method on a novel numerical phantom as well as an experimentally collected phantom and in vivo data. Our code and numerical phantom are available at https://github.com/Lukeli0425/NF-APACT.
Photoacoustic Computed Tomography, Inverse Problems, Neural Fields
1 Introduction
Photoacoustic computed tomography (PACT) images biological samples without cutting through tissue or disrupting normal function. Biological tissues absorb energy from laser stimulation and emit ultrasonic waves, which pass through the rest of the sample and are then collected by an external array of ultrasound transducers. An image of the initial pressure distribution (2D for the scope of this paper) is obtained by processing the ultrasonic signals collected by the transducer array. Traditional image reconstruction algorithms include filtered back-projection [1] and delay-and-sum (DAS) [2, 3, 4], which back-project the collected signals into the image using times of flight and assuming a constant SOS throughout the sample. This assumption neglects the impact of the sample’s internal structure and leads to image aberrations. Specifically, the inherent acoustic heterogeneity, or spatial variations in the 2D SOS distribution, will cause wavefront distortions in the ultrasonic waves that speed or delay their arrival to different parts of the transducer array. Visual artifacts of these distortions include ringing, smear, and doubling of features, which can obscure key sample structures.
A wide range of methods have been developed to mitigate or remove the artifacts caused by acoustic heterogeneity, as recently reviewed in [5, 6]. Autofocus approaches optimize an assumed-constant SOS based on the reconstructed image quality [7] or the coherence factor of the PA signals [8, 9]. Half-time reconstruction methods [10, 11] likewise assume a constant SOS but alleviate aberrations by identifying and discarding regions of the signals that are more affected by acoustic heterogeneity. Dual-SOS DAS [12] efficiently corrects for the first-order effect of acoustic heterogeneity by representing the sample as an oval-shaped region of altered SOS and assuming values for its shape and internally-constant SOS. On the other end of the complexity scale, Ultrasound Computed Tomography (USCT)-enhanced methods [13, 14, 15] directly measure the 2D SOS with a complementary ultrasound measurement apparatus. This measurement enhances image reconstruction but is challenging and expensive to implement simultaneously with PACT and USCT measurement procedures at the same time. The methods mentioned above either require oversimple assumptions or costly measurement of the SOS, and can only unreliably mitigate instead of fully undoing the SOS aberration in practice, see [5].
In contrast to methods that assume or measure SOS, joint reconstruction approaches simultaneously recover images and 2D SOS from PA signals. Regularization-based joint reconstruction [17, 18] alternatively reconstructs the image and SOS in an iterative framework with regularization on both distributions. However, the ill-posed nature of the problem causes numerical instability and error accumulation. The feature coupling method [19] parameterizes the SOS with several constant value regions and maximizes the similarity between reconstructed images from partial arrays. Its drawbacks are the inability to model a complicated SOS map and reliance on a good initial guess. Inspired by indirect wavefront sensing in adaptive optics, Adaptive PACT (APACT) [20] divides the image into small patches and solves for the wavefront in every patch via an exhaustive search. APACT produces state-of-the-art image quality, but with two major drawbacks. First, APACT assumes a simplified wavefront model that cannot account for overall scaling of the reconstructed image. Second, its exhaustive search is computationally expensive, leading to extended reconstruction times that limit its practical application. Our joint reconstruction approach addresses these weaknesses with a novel deep learning approach.
Deep learning has been applied to SOS aberration correction [21, 22] and other image restoration tasks [23] in PACT. However, a key challenge lies in generalizing models trained on large simulated datasets to real-world in vivo data [6]. The scarcity of realistic, high-fidelity simulated data causes a domain shift between training and testing that can significantly degrade results on real data. We see a way around this challenge in the ability of neural networks to represent sample properties as functions of spatial coordinates, learned directly from observations in a self-supervised manner [16, 24]. These methods, termed implicit neural representations (INR) or neural fields (NF), have shown remarkable success in novel view synthesis [25, 26, 27], medical image reconstruction [28, 29, 30, 31, 32, 33], intensity diffraction tomography [34], and Fourier ptychographic microscopy [35]. NFs train a relatively small number of parameters with no need for external or labeled training data. This property is instrumental in scientific applications, where large labeled datasets are scarce.
Inspired by the success of NFs and APACT, we present Neural Fields for Adaptive Photoacoustic Computed Tomography (NF-APACT), a self-supervised framework for joint reconstruction of 2D SOS and initial pressure images in PACT, illustrated in Fig. 1. Our key contributions are:
1. Our method achieves state-of-the-art image quality in a fraction of the compute time ( faster), while simultaneously recovering an approximate SOS. It is robust to nonideal effects (e.g. transducer EIRs and geometric offsets) found in real PACT systems, as demonstrated by its translation to in vivo data without modification.
2. We use no external training data, instead using only PA signals for self-supervised test-time training. Our method learns an SOS that reproduces the observed image aberrations by backpropagation through efficient and differentiable forward and inverse modules. The physics-based forward module incorporates a parameterization of the SOS by a neural field that reduces the number of unknowns by 2 orders of magnitude for fast and stable reconstruction.
3. NF-APACT is an interpretable method that explicitly reports the reconstructed SOS, applies a physics-based forward model from SOS to aberrations, and uses an analytical robust deconvolution for image reconstruction. Domain knowledge can be explicitly incorporated through user-defined regularization of the SOS.
We benchmark the proposed method and previous methods on image quality, compute time on both simulated and realistic data, showing that the proposed method performs better and faster compared to state-of-the-art method APACT [20]. Our code and numerical phantom are available in our GitHub repository (https://github.com/Lukeli0425/NF-APACT).
2 Method
2.1 Image Formation and Aberration
Filtered back-projection [1] is a widely used image reconstruction method in ring array PACT systems. In practice, due to the limited bandwidth and non-flat angular response of the transducers [20], a delay-and-sum (DAS) algorithm [2, 3, 4] is often used as a reasonable approximation. A delay-and-sum image varies over sample locations as well as a user-selected extra delay parameter that functions as a “focus setting” [7], and is reconstructed according to
where is the assumed constant SOS, is the PA signal at time for the n transducer, is the location of the n transducer, and is the total number of transducers.
If the SOS throughout the sample is constant and matches the background, there is a single velocity and delay that will create a sharp image. Because this is not the case for real samples, points in the object are smeared across the image into irregular patterns [36]. The delay value is a mitigation for these aberrations, and can improve image quality bending each back-projected arc, with the risk that poorly-selected delays will spread sample points into rings on the image (see PSFs in first column of Fig. 2). In real samples, with heterogeneous SOS, more complex aberrations arise and cannot be resolved with a single delay parameter (compare to right columns of Fig. 2). Specifically, assuming a straight acoustic rays model [36], the time-of-flight of a signal from sample point to a transducer at is
where is the unknown 2D SOS. From [20], the angle-varying displacement between the actual, distorted wavefront at location for SOS and the uniform wavefront assumed by DAS is the wavefront error :
where is the angle of the vector . See Fig. 2b for example wavefront errors corresponding to the SOS maps shown in panel a for the location indicated by white crosses. It has been shown in [20] that the resulting aberrations, despite varying across the sample, can be modeled within isoplanatic image patches as convolutions with PSFs and unaberrated patches :
Specifically, these PSFs can be computed in the Fourier domain (indicated with capital letters) from the wavefront-based transfer function as
where is the frequency coordinate and is the center location of the i image patch. See Fig. 3a for an illustration of PSFs at varying locations and delays in our numerical phantom. At each location, the delay that generates the most compact PSF is indicated in red.
2.2 Multi-channel Deconvolution
Joint processing across multiple delay values provides more robust deconvolution results. A DAS reconstruction for each extra delay distance comprises our DAS stack, and for each patch of the image stack these delay values (in conjunction with SOS , on which more later) generates a corresponding PSF stack. Because each PSF takes the same underlying clean image into the aberrated image stack, the problem is overdetermined and can be expressed in the frequency domain as
Expressed in matrix form as , we reconstruct clean image with a pseudo-inverse,
This simple method succeeds thanks to several sources of robustness: the combination of information encoded across multiple delay channels, the fact that the digital creation of new channels generates no additional measurement error, and later in our pipeline the consistency across overlapping image patches through Gaussian-weighted merging.
Fig. 3b shows the image patch recovered using single-channel deconvolution with each delay’s PSF. The most compact PSF (red) produces the best result of these, but artifacts are extreme for each individual channel. By using 32 delays for multichannel deconvolution, our method reduces ringing and includes features that are otherwise lost at patch edges, see arrows in Fig. 3c. The reduction of aberrations as more delay channels are added can be seen in Fig. 3d, which also shows a growth in compute time from 0.4 to 1.1 minutes for 4 to 32 channels. While each multichannel deconvolution takes 5.0 seconds, this time is spent iteratively refining our estimate of the SOS, as described in the following sections.
2.3 Implicit Neural SOS Representation
To tackle the challenging problem of SOS reconstruction, we used a non-pretrained, coordinate-based, fully-connected network to parameterize the SOS as a function of sample location:
where are trained network parameters and is a set of spatial coordinates in the SOS map with the same resolution as the recovered image. To reduce computation, we restrict these coordinates to a circular mask roughly the size of the biological sample and set the SOS outside as a constant. This constant background SOS is calculated from the water temperature [37] and is also the SOS used in DAS.
The advantages of this approach are multi-fold. First, implicit neural representations can be learned in a self-supervised manner. This means the network is trained at test time for consistency on the input PA signals alone, which is vital due to the lack of datasets with ground truth images in this domain. Second, by learning a continuous function of the coordinates, the network serves as a natural prior that promotes continuity in the SOS map. Finally, the number of trained parameters is much smaller than the number of pixels to reconstruct in the SOS map. This significantly reduces the degrees of freedom in the inverse problem and makes the problem better-posed.
We used the same network architecture as SIREN [24], where sinusoidal activation functions are utilized to enhance the network’s ability to fit complex signals. Our MLP network consists of one hidden layer with 64 neurons and an output layer. The number of learnable parameters of our network is 4417, which is much smaller than the number of unknown values in the SOS map (200,000). Similar results were produced by networks with different neuron and layer counts, with 1.2-132k trainable parameters.
2.4 Joint Reconstruction
By combining our neural representation of SOS with differentiable PSF generation and deconvolution modules, shown in red in Fig. 1, we are able to backpropagate through the entire forward and inverse models to achieve consistency between the expected and observed image aberrations. This allows joint reconstruction of the unaberrated image and the SOS from PA signals with no external training data. Specifically, we train the SOS network parameters to minimize the loss
over all image patches simultaneously. The first of these terms imposes a physics-informed data fidelity that includes both the forward model in as well as a Fourier-noise-specific weighting as in [20]. The second term smooths the SOS directly with total variation regularization [39]. Training is performed only on the SOS representation parameters but requiring accurate aberration reconstruction across all delays and image patches leads to sufficiently accurate PSFs that the final multichannel deconvolution produces high-quality recovered image patches . Finally, these patches are merged into a single image by applying a Gaussian window with a FWHM of 1.5 mm to the image patches [20] for smooth interpatch transitions and adding the patches back to the initial pressure at corresponding patch locations.
We used Adam optimizer with a learning rate of and trained for 10 epochs. We select the learning rate and epoch count that empirically led to convergence in shortest compute time. Our method is implemented using PyTorch and is available at https://github.com/Lukeli0425/NF-APACT.
3 Experiments
3.1 Numerical Simulations
To test the proposed method with a ground truth reference, we designed a numerical phantom based on mouse liver structure, shown in Fig. 4a. The background water SOS is set to 1499.4 m/s (assuming a 26∘C water temperature based on [37]) outside the masked area, and SOS ranged from 1510.0 to 1650.0 m/s. The PA signals are obtained by 2D numerical simulation using the k-Wave toolbox [38] assuming an evenly distributed 512-transducer ring array with a diameter of 10 cm (same setting for the following phantom and in vivo experiments). Similar to [20], we take the derivative of the PA signals to simulate transducer effects: . The images are cropped into 75% overlapping 3.2mm 3.2mm patches, as for all reported results.
The wavefront errors modeled/reconstructed (sample location marked by white “x” on SOS in Fig. 4a) by different methods are shown in Fig. 4b. To use the terminology of [20], the wavefront errors can be decomposed into Fourier series, where the components degrade the image differently. Conventional DAS assumes a uniform SOS thus a constant wavefront (yellow), while Dual-SOS DAS assumes an additional body SOS and is able to capture the zeroth and first-order term accurately (green). APACT includes only the zeroth and second order Fourier series terms in its wavefront model (blue), missing both the first order (the cosine component captured by dual-SOS in green) as well as higher order terms capturing finer detail. In comparison, NF-APACT can accurately model lower and higher-order components in the wavefront (purple), compared with the truth wavefront (red).
These wavefront errors are directly reflected in image reconstruction quality in Fig. 4c. Conventional DAS, which assumes a uniform SOS of 1510.0 m/s, not only suffers from doubling artifacts (see insets) due to the ignorance of higher order terms but also causes the image to shrink (see vertical shift of the circle and boundary in green box). This shrinking is caused by the missing first order term in the wavefront, which leads to off-center PSFs that shift the image patches toward the center of the tissue. Note also that for real samples this uniform SOS value is tuned by generating many reconstructions and requiring the user to select the best image, which significantly increases the reconstruction time in practice. Dual-SOS DAS assumes a body SOS of 1561.0 m/s (also hand-tuned with search time not included in reported reconstruction time) and is able to fix the shrinking problem but artifacts still remain. The image reconstructed by APACT is also shrunk (see green box and the image quality metrics) due to ignorance of the first-order term. APACT oversmooths the SOS and loses key features in the SOS, e.g., the orange liver-like region and the hole in the upper right, and is also slow and expensive due to the exhaustive search for wavefronts. Our method reconstructs a more accurate SOS, corresponding to a significant improvement in image quality over traditional methods. Image patches show fewer aberrations than all previous methods, reflected in high PSNR and SSIM scores. Our reconstruction takes only 1.1 minutes, compared to 37 minutes for APACT. We also show the image reconstructed by multi-channel deconvolution using the ground truth SOS as an upper limit baseline for image quality and speed of USCT-enhanced reconstruction.
3.2 Leaf Phantom
We use the phantom data from [20], which features a leaf in a 2.5 cm diameter agarose cylinder with a water-filled hole in the middle for acoustic heterogeneity (see photograph in lower left). The phantom is surrounded by water with a temperature of 26∘C and the wavelength of laser illumination is 700 nm in the experiment. Similarly, the constant background SOS is set to 1499.4 m/s according to [37]. Due to nonideal transducer alignment and transducer impulse responses, the collected PA signals are calibrated using measured ring error and transducer electrical impulse response (EIR) before feeding into the algorithms. The same calibration is also applied to the following in vivo experiment.
The reconstruction results of NF-APACT and several other methods are shown in Fig. 5a. The lower left of the phantom is dim due to the nonuniform illumination [20]. We searched over SOS values to report the best outputs, using SOS values of 1505.0 m/s for conventional DAS and 1520.0 m/s for body SOS in Dual-SOS DAS. NF-APACT is able to refocus the leaf veins in the cylinder (blue) as well as inside the hole (red), where the signals are weak due to nonuniform and acoustic reflections, and DAS and APACT both fail. The SOS reconstructed by NF-APACT has a lower SOS region in the middle matching the hole filled with water, while APACT smooths out the SOS and cannot resolve the hole. In terms of compute time, our method is also an order of magnitude faster than APACT.
3.3 In vivo Mouse Liver
In our in vivo experiment, we use the PA signals collected from a nude mouse liver at 1064 nm in [19]. An illustration of the imaged section is shown in the lower left corner. The water temperature was 31∘C during the experiment, thus we used a constant background SOS of 1511.4 m/s. We used 1515.0 m/s for conventional DAS and 1560.0 m/s for Dual-SOS DAS.
The reconstructed results are shown in Fig. 5b. It is hard to fully characterize the reconstruction qualities without ground truth, but we are able to demonstrate the success of our method by looking at several important features. NF-APACT is able to increase the overall contrast and recover structures like blood vessels (green and blue) and bright bone structures (red). Our reconstructed SOS has a superior match to the liver’s anatomical shape. In comparison, APACT reconstructs a smooth SOS in which the middle, high-SOS region does not align well with the liver structure.
4 Discussion and Conclusion
We have introduced NF-APACT, a self-supervised framework achieving fast, state-of-the-art performance in joint reconstruction of initial pressure and SOS in PACT. This method combines the neural field representation from computer graphics with an extension of the optical metaphors of autofocus and adaptive optics used in previous methods to a multichannel image representation analogous to a focal stack. We highlight the following as the key features of our novel method:
Reliability and interpretability. NF-APACT combines signal processing with neural networks to provide reasonable guess of the SOS without a black box. In contrast to deep learning methods where neural networks are trained on large datasets, our method utilizes a small MLP as an efficient representation in a physics-based framework. Our approach enhances reliability and interpretability, making it well-suited for scientific and medical applications where black-box models are often undesirable.
Time-performance trade-offs. Existing reconstruction methods offer diverse solutions for different demands. Based on our benchmarking, we can recommend Dual-SOS DAS with a properly-selected SOS for real-time removal of many aberrations. When image quality is key, we offer a new method to access state-of-the-art reconstruction of initial pressure and SOS in roughly 1 minute. The number of training epochs and learning rate in NF-APACT further allow for customization of the time-performance trade-off. Additionally, we demonstrate 6-second reconstruction when SOS is measured directly.
Generalizability. Although only demonstrated in 2D, our framework could potentially extend to 3D PACT systems. In 3D PACT, where 2D scans from various locations and orientations are combined to reconstruct a 3D volume, the major challenge lies in the large number of unknowns in the 3D SOS, which scales up the computational cost by order of magnitudes. To address this challenge, the efficient SOS representation combined with the efficient and differentiable forward model in NF-APACT offers a promising solution. By reducing the degree of freedoms and lowering computational costs, NF-APACT could enable practical 3D PACT applications.
Acknowledgment
We would like to thank Liujie Gu and Yan Luo for helping us with the numerical simulations.

### v3

@@
-Neural Fields for Adaptive Photoacoustic Computed Tomography
+Self-Supervised Speed of Sound Recovery for Aberration-Corrected Photoacoustic Computed Tomography
 Abstract
-Photoacoustic computed tomography (PACT) is a non-invasive imaging modality with wide medical applications.
-Conventional PACT image reconstruction algorithms suffer from wavefront distortion caused by the heterogeneous speed of sound (SOS) in tissue, which leads to image degradation.
-Accounting for these effects improves image quality, but measuring the SOS distribution is experimentally expensive.
-An alternative approach is to perform joint reconstruction of the initial pressure image and SOS using only the PA signals.
-Existing joint reconstruction methods come with limitations: high computational cost, inability to directly recover SOS, and reliance on inaccurate simplifying assumptions.
-Implicit neural representation, or neural fields, is an emerging technique in computer vision to learn an efficient and continuous representation of physical fields with a coordinate-based neural network.
-In this work, we introduce NF-APACT, an efficient self-supervised framework utilizing neural fields to estimate the SOS in service of an accurate and robust multi-channel deconvolution.
-Our method removes SOS aberrations an order of magnitude faster and more accurately than existing methods.
-We demonstrate the success of our method on a novel numerical phantom as well as an experimentally collected phantom and in vivo data.
-Our code and numerical phantom are available at https://github.com/Lukeli0425/NF-APACT.
-Photoacoustic Computed Tomography, Inverse Problems, Neural Fields
+Photoacoustic computed tomography (PACT) is a non-invasive imaging modality, similar to ultrasound, with wide-ranging medical applications.
+Conventional PACT images are degraded by wavefront distortion caused by the heterogeneous speed of sound (SOS) in tissue.
+Accounting for these effects can improve image quality and provide medically useful information, but measuring the SOS directly is burdensome and the existing joint reconstruction method is computationally expensive.
+Traditional supervised learning techniques are currently inaccessible in this data-starved domain.
+In this work, we introduce an efficient, self-supervised joint reconstruction method that recovers SOS and high-quality images using a differentiable physics model to solve the semi-blind inverse problem.
+The SOS, parametrized by either a pixel grid or a neural field (NF), is updated directly by backpropagation.
+Our method removes SOS aberrations more accurately and 35x faster than the current SOTA.
+We demonstrate the success of our method quantitatively in simulation and qualitatively on experimentally-collected and in-vivo data.
 1 Introduction
-Photoacoustic computed tomography (PACT) images biological samples without cutting through tissue or disrupting normal function.
-Biological tissues absorb energy from laser stimulation and emit ultrasonic waves, which pass through the rest of the sample and are then collected by an external array of ultrasound transducers.
-An image of the initial pressure distribution (2D for the scope of this paper) is obtained by processing the ultrasonic signals collected by the transducer array.
-Traditional image reconstruction algorithms include filtered back-projection [1] and delay-and-sum (DAS) [2, 3, 4], which back-project the collected signals into the image using times of flight and assuming a constant SOS throughout the sample.
+Photoacoustic computed tomography (PACT) images biological samples without cutting through tissue or disrupting normal function [36].
+Biological tissues absorb energy from laser stimulation and emit ultrasonic waves, which propagate through the rest of the sample and are then collected by an external array of ultrasound transducers.
+An image of the initial pressure (IP) distribution (2D for the scope of this paper) is obtained by processing the ultrasonic signals collected by the transducer array.
+Traditional image reconstruction algorithms include filtered back-projection [40] and delay-and-sum (DAS) [31, 10, 12], which back-project the collected signals into the image using times of flight assuming a constant SOS throughout the sample.
 This assumption neglects the impact of the sample’s internal structure and leads to image aberrations.
 Specifically, the inherent acoustic heterogeneity, or spatial variations in the 2D SOS distribution, will cause wavefront distortions in the ultrasonic waves that speed or delay their arrival to different parts of the transducer array.
-Visual artifacts of these distortions include ringing, smear, and doubling of features, which can obscure key sample structures.
-A wide range of methods have been developed to mitigate or remove the artifacts caused by acoustic heterogeneity, as recently reviewed in [5, 6].
-Autofocus approaches optimize an assumed-constant SOS based on the reconstructed image quality [7] or the coherence factor of the PA signals [8, 9].
+Visual artifacts of these distortions include ringing, smear, and doubling of features, which can obscure key sample structures [41].
+Though supervised learning methods have made significant progress on image restoration in many applications, PACT is a relatively novel imaging modality and appropriate datasets (large and naturally distributed set of measurements, with realistic aberrations, accompanied by ground truth for both image and SOS values) are not available.
+We address these challenges with a self-supervised joint recovery method that learns SOS by backpropagation through a differentiable physical model, then uses the SOS to reconstruct a clean image.
+Specifically, as shown in in Fig.
+1, we optimize a differentiable representation of the SOS, parameterized by a pixel grid (PG) or neural field (NF), to best match observed aberrations, which can then be removed with deconvolution.
+By imposing physical consistency across image locations and a delay-based “focal stack” (details in Sec.
+3), this method converges quickly to produce an approximate SOS and high quality image.
+Our key contributions are:
+1. A physics-based and interpretable PACT imaging model that is fully differentiable, enabling time- and memory-efficient optimization through backpropagation.
+2. A more accurate, globally consistent representation of wavefront errors, for SOTA reconstruction quality in both image and SOS.
+3. A self-supervised, test-time training framework that requires no external training data and can flexibly incorporate novel SOS representations and regularizers.
+We benchmark the proposed method and previous methods on simulated data, showing that our method achieves state-of-the-art reconstructions in a fraction of the compute time (40 sec vs.
+23 min).
+In qualitative results on real data, we show the method is robust to nonideal effects (e.g.
+transducer responses and geometric offsets) found in PACT systems, as demonstrated by its translation to in vivo data without modification.
+2 Related Work
+Direct SOS Aberration Mitigation
+A wide range of methods have been developed to mitigate or remove the artifacts caused by acoustic heterogeneity in PACT, as recently reviewed in [34, 30].
+Autofocus approaches optimize an assumed-constant SOS based on the reconstructed image quality [33] or the coherence factor of the photoacoustic (PA) signals [43, 5].
 Half-time reconstruction methods [1, 23] likewise assume a constant SOS but alleviate aberrations by identifying and discarding regions of the signals that are more affected by acoustic heterogeneity.
-Dual-SOS DAS [12] efficiently corrects for the first-order effect of acoustic heterogeneity by representing the sample as an oval-shaped region of altered SOS and assuming values for its shape and internally-constant SOS.
+Among these methods, Dual-SOS DAS [18] provides the best image reconstruction quality, by efficiently correcting the first-order effect of acoustic heterogeneity by representing the sample with an oval-shaped region of altered SOS and assuming values for its shape and internally-constant SOS.
 On the other end of the complexity scale, Ultrasound Computed Tomography (USCT)-enhanced methods [15, 19, 35] directly measure the 2D SOS with a complementary ultrasound measurement apparatus.
-This measurement enhances image reconstruction but is challenging and expensive to implement simultaneously with PACT and USCT measurement procedures at the same time.
-The methods mentioned above either require oversimple assumptions or costly measurement of the SOS, and can only unreliably mitigate instead of fully undoing the SOS aberration in practice, see [5].
+This measurement enhances image reconstruction but is challenging and expensive to implement simultaneously with PACT.
+The above methods either require oversimple assumptions or costly measurement of the SOS, and can only unreliably mitigate instead of fully undoing the SOS aberration in practice [34].
+Joint Recovery of SOS and Image
 In contrast to methods that assume or measure SOS, joint reconstruction approaches simultaneously recover images and 2D SOS from PA signals.
@@
 Its drawbacks are the inability to model a complicated SOS map and reliance on a good initial guess.
-Inspired by indirect wavefront sensing in adaptive optics, Adaptive PACT (APACT) [20] divides the image into small patches and solves for the wavefront in every patch via an exhaustive search.
-APACT produces state-of-the-art image quality, but with two major drawbacks.
+Inspired by indirect wavefront sensing in adaptive optics (AO) [14, 9, 46], Adaptive PACT (APACT) [6] divides the image into small patches and solves for the wavefront in every patch via an exhaustive search.
+Of these methods, APACT produces state-of-the-art image quality, but with two major drawbacks.
 First, APACT assumes a simplified wavefront model that cannot account for overall scaling of the reconstructed image.
 Second, its exhaustive search is computationally expensive, leading to extended reconstruction times that limit its practical application.
-Our joint reconstruction approach addresses these weaknesses with a novel deep learning approach.
+Our method addresses these weaknesses with a novel self-supervised, physics-based deep learning approach.
+Supervised Learning
 Deep learning has been applied to SOS aberration correction [11, 13] and other image restoration tasks [7] in PACT.
@@
 The scarcity of realistic, high-fidelity simulated data causes a domain shift between training and testing that can significantly degrade results on real data.
-We see a way around this challenge in the ability of neural networks to represent sample properties as functions of spatial coordinates, learned directly from observations in a self-supervised manner [16, 24].
-These methods, termed implicit neural representations (INR) or neural fields (NF), have shown remarkable success in novel view synthesis [25, 26, 27], medical image reconstruction [28, 29, 30, 31, 32, 33], intensity diffraction tomography [34], and Fourier ptychographic microscopy [35].
-NFs train a relatively small number of parameters with no need for external or labeled training data.
-This property is instrumental in scientific applications, where large labeled datasets are scarce.
-Inspired by the success of NFs and APACT, we present Neural Fields for Adaptive Photoacoustic Computed Tomography (NF-APACT), a self-supervised framework for joint reconstruction of 2D SOS and initial pressure images in PACT, illustrated in Fig.
-1.
-Our key contributions are:
-1.
-Our method achieves state-of-the-art image quality in a fraction of the compute time ( faster), while simultaneously recovering an approximate SOS.
-It is robust to nonideal effects (e.g.
-transducer EIRs and geometric offsets) found in real PACT systems, as demonstrated by its translation to in vivo data without modification.
-2.
-We use no external training data, instead using only PA signals for self-supervised test-time training.
-Our method learns an SOS that reproduces the observed image aberrations by backpropagation through efficient and differentiable forward and inverse modules.
-The physics-based forward module incorporates a parameterization of the SOS by a neural field that reduces the number of unknowns by 2 orders of magnitude for fast and stable reconstruction.
-3.
-NF-APACT is an interpretable method that explicitly reports the reconstructed SOS, applies a physics-based forward model from SOS to aberrations, and uses an analytical robust deconvolution for image reconstruction.
-Domain knowledge can be explicitly incorporated through user-defined regularization of the SOS.
-We benchmark the proposed method and previous methods on image quality, compute time on both simulated and realistic data, showing that the proposed method performs better and faster compared to state-of-the-art method APACT [20].
-Our code and numerical phantom are available in our GitHub repository (https://github.com/Lukeli0425/NF-APACT).
+Our physics-based framework, in contrast, shows robust performance across both simulated and in vivo data.
+Neural Fields
+Neural fields (NFs), or implicit neural representations (INRs) have been shown to be capable of representing a variety of natural signals [27, 38], and have been demonstrated in many computational imaging tasks, including medical imaging [21, 24, 44, 29, 26, 39].
+NFs represent images with a relatively small number of parameters compared to pixel grids [28], with no need for external or labeled training data.
+Recent works in PACT have utilized NFs for IP image reconstruction from limited or sparse measurements [47, 37, 42], but these assume constant SOS and only reconstruct the IP image.
+NFs have also been used to reconstruct SOS in ultrasound CT [2] and have shown promise in estimating wavefronts for computational image reconstruction [16] and closed-loop control systems for adaptive optics [8].
+We, like [16], do not intervene in the measurement process, but computationally correct for wavefront-based aberration parameters estimated with a NF.
+We use the SIREN NF architecture, which is a fully connected network with sinusoidal activations [27].
 3 Method
@@
 2).
-Specifically, assuming a straight acoustic rays model [36], the time-of-flight of a signal from sample point to a transducer at is
+Specifically, assuming a straight acoustic ray model [41], the time-of-flight of a signal from sample point to a transducer at is
 where is the unknown 2D SOS. From [6], the angle-varying displacement between the actual, distorted wavefront at location for SOS and the uniform wavefront assumed by DAS is the wavefront error :
@@
 Joint processing across multiple delay values provides more robust deconvolution results.
-A DAS reconstruction for each extra delay distance comprises our DAS stack, and for each patch of the image stack these delay values (in conjunction with SOS , on which more later) generates a corresponding PSF stack.
+A DAS reconstruction for each extra delay distance comprises our DAS stack, and for each patch of this image stack these delay values (in conjunction with SOS , on which more later) generate a corresponding PSF stack.
 Because each PSF takes the same underlying clean image into the aberrated image stack, the problem is overdetermined and can be expressed in the frequency domain as
-Expressed in matrix form as , we reconstruct clean image with a pseudo-inverse,
+and in matrix form as . We reconstruct clean image patches with a pseudo-inverse,
 This simple method succeeds thanks to several sources of robustness: the combination of information encoded across multiple delay channels, the fact that the digital creation of new channels generates no additional measurement error, and later in our pipeline the consistency across overlapping image patches through Gaussian-weighted merging.
@@
 The most compact PSF (red) produces the best result of these, but artifacts are extreme for each individual channel.
-By using 32 delays for multichannel deconvolution, our method reduces ringing and includes features that are otherwise lost at patch edges, see arrows in Fig.
+By using 16 delays for multichannel deconvolution, our method reduces ringing and includes features that are otherwise lost at patch edges, see arrows in Fig.
 3c.
 The reduction of aberrations as more delay channels are added can be seen in Fig.
-3d, which also shows a growth in compute time from 0.4 to 1.1 minutes for 4 to 32 channels.
-While each multichannel deconvolution takes 5.0 seconds, this time is spent iteratively refining our estimate of the SOS, as described in the following sections.
-2.3 Implicit Neural SOS Representation
-To tackle the challenging problem of SOS reconstruction, we used a non-pretrained, coordinate-based, fully-connected network to parameterize the SOS as a function of sample location:
-where are trained network parameters and is a set of spatial coordinates in the SOS map with the same resolution as the recovered image.
-To reduce computation, we restrict these coordinates to a circular mask roughly the size of the biological sample and set the SOS outside as a constant.
-This constant background SOS is calculated from the water temperature [37] and is also the SOS used in DAS.
-The advantages of this approach are multi-fold.
-First, implicit neural representations can be learned in a self-supervised manner.
-This means the network is trained at test time for consistency on the input PA signals alone, which is vital due to the lack of datasets with ground truth images in this domain.
-Second, by learning a continuous function of the coordinates, the network serves as a natural prior that promotes continuity in the SOS map.
-Finally, the number of trained parameters is much smaller than the number of pixels to reconstruct in the SOS map.
-This significantly reduces the degrees of freedom in the inverse problem and makes the problem better-posed.
-We used the same network architecture as SIREN [24], where sinusoidal activation functions are utilized to enhance the network’s ability to fit complex signals.
-Our MLP network consists of one hidden layer with 64 neurons and an output layer.
-The number of learnable parameters of our network is 4417, which is much smaller than the number of unknown values in the SOS map (200,000).
-Similar results were produced by networks with different neuron and layer counts, with 1.2-132k trainable parameters.
+3d, which also shows a growth in compute time from 0.9 to 5.0 seconds for 4 to 32 channels.
 3.3 Joint Reconstruction
-By combining our neural representation of SOS with differentiable PSF generation and deconvolution modules, shown in red in Fig.
+By combining our differentiable PSF generation and deconvolution modules, shown in red in Fig.
 1, we are able to backpropagate through the entire forward and inverse models to achieve consistency between the expected and observed image aberrations.
-This allows joint reconstruction of the unaberrated image and the SOS from PA signals with no external training data.
-Specifically, we train the SOS network parameters to minimize the loss
-over all image patches simultaneously.
-The first of these terms imposes a physics-informed data fidelity that includes both the forward model in as well as a Fourier-noise-specific weighting as in [20].
-The second term smooths the SOS directly with total variation regularization [39].
-Training is performed only on the SOS representation parameters but requiring accurate aberration reconstruction across all delays and image patches leads to sufficiently accurate PSFs that the final multichannel deconvolution produces high-quality recovered image patches .
-Finally, these patches are merged into a single image by applying a Gaussian window with a FWHM of 1.5 mm to the image patches [20] for smooth interpatch transitions and adding the patches back to the initial pressure at corresponding patch locations.
-We used Adam optimizer with a learning rate of and trained for 10 epochs.
-We select the learning rate and epoch count that empirically led to convergence in shortest compute time.
-Our method is implemented using PyTorch and is available at https://github.com/Lukeli0425/NF-APACT.
-3 Experiments
+This allows joint reconstruction of the unaberrated image and the SOS from PA signals with no external training data by leveraging the redundant information encoded in the DAS stack.
+We benchmark two types of SOS representation: a pixel grid (PG) with the same resolution as the IP image, and a neural field (NF) representation.
+For the latter, we used a single-layer, 256-feature, fully-connected network with sinusoidal activations [27].
+This non-pretrained, coordinate-based network maps pixel coordinates to SOS predictions.
+The total number of NF parameters (1027) is much smaller than the number of pixels (200k) in the SOS map, which reduces the degrees of freedom in the inverse problem and acts as an implicit regularizer that makes the problem better-posed.
+See Sec.
+S3.3 for ablations on network size.
+For both representations, we reduce unnecessary computation by restricting these pixel grid or NF coordinates to a circular mask roughly the size of the sample and fix the background SOS based on measured water temperature [20]
+We train the SOS parameters to minimize a self-supervised aberrated-image-matching loss:
+over all image patches simultaneously, where is the parameters in the PG or NF.
+This loss function imposes a physics-informed data fidelity that includes both the forward model in as well as a Fourier-noise-specific weighting as in [6].
+We use TV loss to regularize the conventional pixel grid representation, while for the NF representation, we rely instead on the network’s implicit regularization and set .
+Training is performed only on the SOS representation parameters but requiring accurate aberration reconstruction across all delays and image patches leads to sufficiently accurate PSFs that the final multichannel deconvolution produces high-quality recovered IP image patches .
+Finally, these patches are merged into a single image by applying a Gaussian window with a FWHM of 1.5 mm to the image patches [6] for smooth interpatch transitions and adding the patches back to the IP image at corresponding patch locations.
+We use a delay count of by default in this paper (see ablations in Sec.
+4.4).
+See Sec.
+S1.4 for details on our image patches and stitching.
+4 Results
 4.1 Numerical Simulations
-To test the proposed method with a ground truth reference, we designed a numerical phantom based on mouse liver structure, shown in Fig.
-4a.
-The background water SOS is set to 1499.4 m/s (assuming a 26∘C water temperature based on [37]) outside the masked area, and SOS ranged from 1510.0 to 1650.0 m/s.
-The PA signals are obtained by 2D numerical simulation using the k-Wave toolbox [38] assuming an evenly distributed 512-transducer ring array with a diameter of 10 cm (same setting for the following phantom and in vivo experiments).
-Similar to [20], we take the derivative of the PA signals to simulate transducer effects: .
-The images are cropped into 75% overlapping 3.2mm 3.2mm patches, as for all reported results.
+To quantitatively benchmark the performance of our method with ground truth references, we designed a set of numerical phantoms with 5 different SOS maps (see Sec.
+S1.1 for details), one of which is shown in Fig.
+4a for visualization of the results.
 The wavefront errors modeled/reconstructed (sample location marked by white “x” on SOS in Fig.
@@
 4b.
-To use the terminology of [20], the wavefront errors can be decomposed into Fourier series, where the components degrade the image differently.
-Conventional DAS assumes a uniform SOS thus a constant wavefront (yellow), while Dual-SOS DAS assumes an additional body SOS and is able to capture the zeroth and first-order term accurately (green).
-APACT includes only the zeroth and second order Fourier series terms in its wavefront model (blue), missing both the first order (the cosine component captured by dual-SOS in green) as well as higher order terms capturing finer detail.
-In comparison, NF-APACT can accurately model lower and higher-order components in the wavefront (purple), compared with the truth wavefront (red).
+To use the terminology of [6], the wavefront errors can be decomposed into a Fourier series, where the components degrade the IP image differently.
+Conventional DAS assumes a uniform SOS thus a constant wavefront (red), while Dual-SOS DAS assumes an additional body SOS and is able to capture the zeroth and first-order term accurately (yellow).
+APACT (green) includes only the zeroth and second order Fourier series terms in its wavefront model (see [6] and Sec.
+S2 for details), missing both the first order (the cosine component captured by dual-SOS in yellow) as well as higher order terms capturing finer detail.
+In comparison, our methods can accurately model lower and higher-order components in the wavefront (purple and blue), compared with the truth wavefront (black).
 These wavefront errors are directly reflected in image reconstruction quality in Fig.
@@
 Conventional DAS, which assumes a uniform SOS of 1510.0 m/s, not only suffers from doubling artifacts (see insets) due to the ignorance of higher order terms but also causes the image to shrink (see vertical shift of the circle and boundary in green box).
-This shrinking is caused by the missing first order term in the wavefront, which leads to off-center PSFs that shift the image patches toward the center of the tissue.
-Note also that for real samples this uniform SOS value is tuned by generating many reconstructions and requiring the user to select the best image, which significantly increases the reconstruction time in practice.
-Dual-SOS DAS assumes a body SOS of 1561.0 m/s (also hand-tuned with search time not included in reported reconstruction time) and is able to fix the shrinking problem but artifacts still remain.
+This shrinking is caused by the missing first-order term in the wavefront, which leads to off-center PSFs that shift the image patches toward the center of the tissue.
+Note also that for real samples this uniform SOS value is tuned by generating many reconstructions and requiring the user to select the best image, significantly increasing the reconstruction time in practice.
+Dual-SOS DAS assumes a body SOS of 1561.0 m/s (also hand-tuned with search time not included in reported time) and is able to fix the shrinking problem but artifacts remain.
 The image reconstructed by APACT is also shrunk (see green box and the image quality metrics) due to ignorance of the first-order term.
-APACT oversmooths the SOS and loses key features in the SOS, e.g., the orange liver-like region and the hole in the upper right, and is also slow and expensive due to the exhaustive search for wavefronts.
+APACT oversmooths the SOS and loses key features in the SOS, e.g., the orange liver-like region and the hole in the upper right, and is also slow and expensive due to the exhaustive search for wavefronts coefficients.
 Our method reconstructs a more accurate SOS, corresponding to a significant improvement in image quality over traditional methods.
 Image patches show fewer aberrations than all previous methods, reflected in high PSNR and SSIM scores.
-Our reconstruction takes only 1.1 minutes, compared to 37 minutes for APACT.
+Our method with NF takes only 40 seconds, compared to 22 minutes for APACT.
+Our method with PG produces decent IP images in similar compute time but suffers from stripe-like SOS artifacts that originate from the path-integral nature of the wavefront calculation (see Eq.
+3) in the straight acoustic ray model.
 We also show the image reconstructed by multi-channel deconvolution using the ground truth SOS as an upper limit baseline for image quality and speed of USCT-enhanced reconstruction.
+See Tab.
+1 for quantitative evaluation of all the methods, averaged over five numerical phantoms.
+Supplementary videos show the convergence of the IP image and SOS for each experiment.
+| Method |
+Time | ||||||||
+| DAS | 21.49 | 0.372 | - | - | 0.13 s | ||||||||
+| Dual-SOS DAS [18] | 24.42 | 0.446 | - | - | 0.12 s | ||||||||
+| APACT [6] | 21.49 | 0.434 | 17.74 | 0.908 | 23 min | ||||||||
+| Ours (PG) | 25.05 | 0.514 | 21.26 | 0.903 | 113.3 s | ||||||||
+| Ours (NF) | 25.08 | 0.519 | 22.29 | 0.931 | 40.3 s | ||||||||
+| Ours (SOS oracle) | 25.61 | 0.537 | - | - | 2.6 s |
 4.2 Leaf Phantom
-We use the phantom data from [20], which features a leaf in a 2.5 cm diameter agarose cylinder with a water-filled hole in the middle for acoustic heterogeneity (see photograph in lower left).
-The phantom is surrounded by water with a temperature of 26∘C and the wavelength of laser illumination is 700 nm in the experiment.
-Similarly, the constant background SOS is set to 1499.4 m/s according to [37].
-Due to nonideal transducer alignment and transducer impulse responses, the collected PA signals are calibrated using measured ring error and transducer electrical impulse response (EIR) before feeding into the algorithms.
-The same calibration is also applied to the following in vivo experiment.
-The reconstruction results of NF-APACT and several other methods are shown in Fig.
+We use the phantom data from [6], which features a leaf in a 2.5 cm diameter agarose cylinder with a water-filled hole in the middle for acoustic heterogeneity (see photograph in lower left of Fig.
+5a and details in Sec.
+S1.2).
+The reconstructions of our methods and the baselines are shown in Fig.
 5a.
-The lower left of the phantom is dim due to the nonuniform illumination [20].
+The lower left of the phantom is dim due to nonuniform illumination [6].
 We searched over SOS values to report the best outputs, using SOS values of 1505.0 m/s for conventional DAS and 1520.0 m/s for body SOS in Dual-SOS DAS.
-NF-APACT is able to refocus the leaf veins in the cylinder (blue) as well as inside the hole (red), where the signals are weak due to nonuniform and acoustic reflections, and DAS and APACT both fail.
-The SOS reconstructed by NF-APACT has a lower SOS region in the middle matching the hole filled with water, while APACT smooths out the SOS and cannot resolve the hole.
-In terms of compute time, our method is also an order of magnitude faster than APACT.
+Ours are able to refocus the leaf veins in the cylinder (blue) as well as inside the hole (red), where the signals are weak and DAS and APACT both fail.
+The SOS reconstructed by our methods have a lower SOS region in the middle matching the hole filled with water, while APACT smooths out the SOS and cannot resolve the hole.
+PG still suffers from stripe-like artifacts in the SOS (zoom in) but also reconstructs a well-resolved IP image.
 4.3 In vivo Mouse Liver
-In our in vivo experiment, we use the PA signals collected from a nude mouse liver at 1064 nm in [19].
-An illustration of the imaged section is shown in the lower left corner.
-The water temperature was 31∘C during the experiment, thus we used a constant background SOS of 1511.4 m/s.
-We used 1515.0 m/s for conventional DAS and 1560.0 m/s for Dual-SOS DAS.
-The reconstructed results are shown in Fig.
-5b.
-It is hard to fully characterize the reconstruction qualities without ground truth, but we are able to demonstrate the success of our method by looking at several important features.
-NF-APACT is able to increase the overall contrast and recover structures like blood vessels (green and blue) and bright bone structures (red).
-Our reconstructed SOS has a superior match to the liver’s anatomical shape.
-In comparison, APACT reconstructs a smooth SOS in which the middle, high-SOS region does not align well with the liver structure.
+In our in vivo experiment, we used a nude mouse liver from [3] (details in Sec.
+S1.3).
+The reconstructions are shown in Fig.
+5b along with a cartoon illustration of the imaged section (bottom left corner).
+We hand-tuned SOS assumptions to 1517.0 m/s for conventional DAS and 1560.0 m/s for Dual-SOS DAS.
+Despite lacking ground truth, we observe success on several important features.
+Our method with NF is able to increase the overall contrast and recover structures like blood vessels (green and blue) and bright spine and rib structures (red) in the IP image.
+Our reconstructed SOS has a superior match to the liver’s anatomical shape (see overlaid figure in Fig.
+S1).
+In comparison, APACT reconstructs a smooth SOS in which the middle, high-SOS region does not align well with the liver.
+Ours with PG also suffers from severe artifacts in the SOS.
+4.4 Ablation Studies
+Our ablation studies focus on the effects of the number of delays , TV regularization , and network size.
+We observe an increase in the performance at the cost of compute time as more extra delay distances are used and more redundancy is created in the DAS stack.
+This improvement gradually diminishes beyond a delay count of 16, which we selected for our implementation (see Fig.
+6).
+For NF network size (see Fig.
+S3 and Sec.
+S3.3), our method fails either when the network is too small (NF’s implicit smoothness regularization is too strong) or large (the NF has a stronger ability to learn high-frequency features and has weaker implicit regularization).
+For the TV regularization of PG (see Fig.
+S2 and Sec.
+S3.2), a larger encourages smoother SOS results and reaches optimal performance at a weight of .
 5 Discussion and Conclusion
-We have introduced NF-APACT, a self-supervised framework achieving fast, state-of-the-art performance in joint reconstruction of initial pressure and SOS in PACT.
+We have introduced a self-supervised framework achieving fast, state-of-the-art performance in joint reconstruction of initial pressure and speed of sound in PACT.
 This method combines the neural field representation from computer graphics with an extension of the optical metaphors of autofocus and adaptive optics used in previous methods to a multichannel image representation analogous to a focal stack.
@@
 Reliability and interpretability.
-NF-APACT combines signal processing with neural networks to provide reasonable guess of the SOS without a black box.
-In contrast to deep learning methods where neural networks are trained on large datasets, our method utilizes a small MLP as an efficient representation in a physics-based framework.
+Our method combines signal processing with neural networks to provide reasonable guess of the SOS without a black box.
+In contrast to deep learning methods where neural networks are trained on large datasets, our method utilizes a small MLP as an efficient representation in a physics-based framework and incorporates redundant information in the measurements to efficiently solve a partially-blind inverse problem.
 Our approach enhances reliability and interpretability, making it well-suited for scientific and medical applications where black-box models are often undesirable.
@@
 Based on our benchmarking, we can recommend Dual-SOS DAS with a properly-selected SOS for real-time removal of many aberrations.
-When image quality is key, we offer a new method to access state-of-the-art reconstruction of initial pressure and SOS in roughly 1 minute.
-The number of training epochs and learning rate in NF-APACT further allow for customization of the time-performance trade-off.
-Additionally, we demonstrate 6-second reconstruction when SOS is measured directly.
+When image quality is key, we offer a new method to access state-of-the-art reconstruction of the IP and SOS in roughly 40 seconds.
+The delay count, network size, and number of training epochs in NF further allow for customization of the time-performance trade-off.
+Additionally, we demonstrate 3-second reconstruction when SOS is measured directly.
 Generalizability.
-Although only demonstrated in 2D, our framework could potentially extend to 3D PACT systems.
-In 3D PACT, where 2D scans from various locations and orientations are combined to reconstruct a 3D volume, the major challenge lies in the large number of unknowns in the 3D SOS, which scales up the computational cost by order of magnitudes.
-To address this challenge, the efficient SOS representation combined with the efficient and differentiable forward model in NF-APACT offers a promising solution.
-By reducing the degree of freedoms and lowering computational costs, NF-APACT could enable practical 3D PACT applications.
-Acknowledgment
-We would like to thank Liujie Gu and Yan Luo for helping us with the numerical simulations.
+Although only demonstrated in 2D, our framework could potentially extend to 3D PACT systems [4].
+In 3D PACT, where multiple 2D scans are combined to reconstruct a 3D volume, the major challenge lies in the large number of unknowns in the 3D SOS, which drastically scales up the computational time and memory requirements.
+To address this challenge, our efficient NF SOS representation combined with efficient and differentiable forward model offers a promising solution.
+By reducing the degrees of freedom and lowering computational costs, the proposed method could enable practical 3D PACT applications.
+Supplementary Material
+S1 Experiment Details
+S1.1 Numerical Phantoms
+In this section, we describe the simulation details of our numerical phantoms.
+We use the same IP image shown in Fig.
+4a for all numerical phantoms but use five SOS with different variations (see Fig.
+S4a).
+SOS values range from 1490.0 to 1650.0 m/s in our numerical dataset.
+The structures in the IP and SOS images are inspired by the mouse body111https://www.imaios.com/en/vet-anatomy/mouse/mouse-whole-body while the SOS values are based on realistic measurements222https://itis.swiss/virtual-population/tissue-properties/database/acoustic-properties/.
+For example, the small circles with high SOS in samples 1 and 3 correspond to bones, the orange circles in samples 1 and 2 correspond to the liver, and the circle with small SOS in samples 1, 4, and 5 correspond to the stomach lumen.
+The background water SOS is set to 1499.4 m/s (assuming a 26∘C water temperature based on [20]) outside the sample.
+The PA signals are obtained by 2D numerical simulation using the k-Wave toolbox [32] assuming an evenly distributed 512-transducer ring array with a diameter of 10 cm (same setting for the leaf phantom and in vivo mouse liver experiments).
+Similar to [6], we take the derivative of the PA signals to simulate transducer effects: .
+The images are cropped into 75% overlapping 3.2mm 3.2mm patches, as for all reported results.
+We show the IP and SOS reconstructions on all five numerical phantoms using our methods and APACT in Fig.
+S4 b,c,d.
+Results are qualitatively similar to the example discussed in the text, and numbers reported in tables are averaged over these five phantoms.
+S1.2 Leaf Phantom
+We use the experimentally collected phantom data from [6], which features a leaf in a 2.5 cm diameter agarose cylinder with a water-filled hole in the middle for acoustic heterogeneity (see photograph in Fig.
+5a).
+The leaf phantom is surrounded by water with a temperature of 26∘C, and the wavelength of laser illumination is 700 nm in the experiment.
+The constant background SOS is set to 1499.4 m/s according to [20].
+Due to nonideal transducer alignment and transducer impulse responses, the collected PA signals are calibrated using measured ring error and transducer electrical impulse response (EIR) before feeding into the algorithms.
+The same calibration is applied to the in vivo experiment.
+S1.3 In vivo Mouse Liver
+The PA signals in our in vivo experiments were collected from a nude mouse liver at 1064 nm [3].
+The water temperature was 31∘C during the experiment, thus we used a background SOS of 1511.4 m/s.
+See figure S1 to see the alignment between recovered SOS and IP, which matches better in our method than others.
+S1.4 Image Stitching
+We adopted the same strategy used for APACT [6] in IP reconstruction.
+This process differs from most existing image reconstruction frameworks because the image is reconstructed patch-by-patch and stitched to form the final result.
+Images are cropped using 3.2 mm 3.2 mm Gaussian windows with an FWHM of 1.5 mm before being fed to our algorithm.
+The patch centers are arranged on a Cartesian grid, and adjacent patches overlap by 75%.
+With our proposed method for solving the speed of sound distribution by utilizing information from all patches, we obtain the wavefront correction model for each isoplanatic patch.
+This simultaneously yields the deconvolved, aberration-corrected image patches.
+These patches are then translated to their corresponding center positions and summed.
+The 75% overlap ensures a smooth transition during image stitching.
+Finally, the result is adjusted to remove the modulation pattern due to the Gaussian windows by dividing by the sum of the Gaussian window weights.
+S2 Difference Between Our Wavefront Model and APACT’s Wavefront Model
+One critical contribution of our work is the improved wavefront model compared to the previous work APACT [6].
+As explained in the main text, the wavefront model represents the wavefront advancement or retardation induced by a nonuniform speed‐of‐sound distribution in tissue relative to an ideal spherical wave in a homogeneous medium.
+Mathematically, the wavefront function , which depicts the propagation length difference relative to the spherical wave when received by the transducers, is a function of the propagation direction .
+This wavefront function can be expanded using a Fourier series:
+In APACT, only three low order harmonics were used, namely DC (constant ), , and .
+This is because, firstly, the first order terms and result in a shifted PSF, making them non-identifiable from the reconstructed images when the true IP is unknown.
+Secondly, the authors omitted the higher order terms, i.e.
+and for , to avoid the exponential increase in computation time during the exhaustive search for the optimal wavefront.
+However, as shown in Fig.
+2, the actual wavefront function can be complex, which necessitate those high order terms.
+In our work, the wavefront functions corresponding to different isoplanatic patches are obtained from the learned SOS map.
+There is no explicit restriction on their profiles such as the assumption in APACT that they are slowly varying.
+S3 Ablation Studies
+We conducted ablation studies using our synthetic numerical phantoms on an NVIDIA RTX A6000 GPU. All results are averaged over 5 numerical phantoms, which share the same IP but have different SOS.
+S3.1 Number of Delays
+We show the performance of our method (with and without SOS oracle), pixel grid (with and without TV regularization), and APACT with respect to numbers of delays used in Fig.
+6.
+We keep the range of the delay distances fixed at [-0.8 mm, 0.8 mm] and uniformly sample delay distances within this range based on the number of delays used.
+We show that more delay channels leads to increased performance and computation time for all methods.
+This originates from Eq.
+7, where more delay channels promote the performance of pseudo-inverse deconvolution, provide more redundant information for SOS estimation, and require more computation time.
+Removing TV regularization in pixel grid (dashed green line) has little impact on the quality of the IP image but significantly decreases the quality of the SOS.
+Considering the performance and computation time, we choose 16 delays (marked with stars).
+We also show that the efficient parameterization of SOS gives NF a slight time advantage compared to PG (see lower right subplot), which takes more iterations to reach convergence.
+S3.2 TV regularization for PG
+In Fig.
+S2, we show the effects of TV regularization weight with our pixel grid SOS representation.
+We observe that the IP image quality drops when , while the SOS accuracy reaches optimum around .
+Therefore, we select (marked with a green star in Fig.
+S2) for all reconstructions in our implementation of PG.
+S3.3 Network Size of NF
+We compared the performance of our method with NF with different network sizes ranging from 131 to 264,707 parameters, shown in Fig.
+S3.
+We show that the implicit smoothness regularization of the MLP decreases as the network becomes larger, resulting in noisier SOS reconstructions and hence worse IP images.
+This can be observed from the comparison between no hidden layer (solid line) and one hidden layer (dashed line) and the overall decreasing trend of the two curves.
+An increase in network size also leads to a small rise in computation time.
+We select 0 hidden layer and 256 hidden features after an overall consideration of the performance and computation time (marked with a purple star).
+S4 Supplementary Videos
+We provide six videos showing the convergence process of the IP image and SOS for our methods with our numerical phantom (nf_numerical_phantom.mp4, and pg_numerical_phantom.mp4), leaf phantom (nf_leaf_phantom.mp4, and pg_leaf_phantom.mp4), and in vivo mouse liver (nf_in_vivo_mouse_liver.mp4 and pg_in_vivo_mouse_liver.mp4).